Calculate Delta G degree (in J) for 3 moles of ele
Answer #1

\DeltaG0   = -nE0Cell *F

           = -3*0.57*96500

          = -165015J

3. \DeltaG0   = -nE0Cell *F

              = -2*0.28*96500

              = -54040J

4. \DeltaG0   = -nE0Cell *F

                = -2*-0.2*96500

               = 38600J

\DeltaG0    = -RTlnK

38600   = -8.314*298*2.303logK

logK     = -38600/5705.8483

logK    = -6.76

    K     = 10-6.76   =1.74*10-7